# Aligning to Thousands of Preferences via System Message Generalization

### v1

@@ A.1 Instruction sampling and filtering
Nectar anthropic-hh 6,037
lmsys-chat-1m 3,119
sharegpt 1,994
ultrachat 783
evol_instruct 775
flan_v2_niv2 583
flan_v2_cot 253
false_qa 176
flan_v2_p3 171
truthful_qa 54
flan_v2_flan2021 33
@@
A.2 Seed preference creation
We hand-crafted 4 dimensions, 18 sub-dimensions, and 107 preference descriptions. The full list keywords of our seed preference data are in Table 6. We took inspiration from works that delineate query-independent properties of language model responses [44, 36, 21], user intentions in instructions [50], machine personalities [24], moral judgment tasks [46], cultural alignment [3], and fine-grained evaluation criteria of language models [27].
Table 6: Keyword of 107 seed preferences
Dimension Subdimension Preference keywords
@@ A.2 Seed preference creation
Appendix B Analysis of Multifaceted Collection
B.1 Diversity of preferences
To measure the diversity of preferences in Multifaceted Collection, we computed ROUGE-L666pypi.org/project/rouge-score/ (F1 score) similarities between every possible pair of preference descriptions assigned to each instruction. Given that there are three preference sets per instruction and four dimensions defined in each preference set, we computed the metric along the four dimensions, i.e., ($3\choose 2\times 4$ scores are obtained per instruction. As depicted in Figure 4, the average ROUGE-L score across all dimensions is approximately 0.21, with the maximum score reaching 0.25. These results indicate a significant diversity among the preferences, confirming that they do not overlap substantially and are well-distributed within Multifaceted Collection. Example subdimensions and keywords of preference descriptions are displayed in Table 7.
@@ B.1 Diversity of preferences
Appendix C Analysis of Multifaceted Bench
C.1 Data curation
We sampled instructions from five benchmarks and excluded instructions that are either used in Multifaceted Collection or overlapped across the benchmarks, and then sampled from the remaining messages to collect 315 unique instructions. We use GPT-4 Turbo-0125 to generate synthetic multifaceted system messages and their aligned reference answers, similar to how we curated the Multifaceted Collection. For each instruction, we generated three system messages and one reference answer, resulting in a total of 945 sets. In this process, we generate system messages with unseen preferences that are not included in the Multifaceted Collection, thereby reducing the possibility of data contamination. The similarity between Multifaceted Collection and Multifaceted Bench is detailed in section 5.
Lastly, in line with previous works [76, 27, 33, 28] indicating that using a score rubric helps in conducting consistent evaluations, we also develop a score rubric to assess the reference answers. The score rubric evaluates how well the model’s response adheres to each of the four preference dimensions that comprise the system message. We create one rubric per dimension, resulting in four rubrics for each response. Each rubric includes an explanation of the dimension and criteria for scoring from 1 to 5 based on how well the model follows it.
C.2 Train-test similarity
To verify whether Multifaceted Bench indeed contains unseen preferences not present in Multifaceted Collection, we measure the overlap between the system prompts of the two datasets by calculating their Rouge-L scores. As shown in figure 5, the average Rouge-L score is around 0.2, indicating a sufficiently low overlap. Additionally, this score is lower than the similarities among training data as depicted in figure 4. This low level of train-test similarity, compared to previous studies [68, 26], confirms that Multifaceted Bench successfully represents unseen tasks.
Appendix D Hyperparameters and computing resources
D.1 Hyperparameters
To train Janus 7B, we utilize the axolotl library777github.com/OpenAccess-AI-Collective/axolotl. For instruction tuning, the configuration includes a maximum sequence length of 8192, gradient accumulation steps of 4, a micro batch size of 2, and four epochs. We use the adamw_bnb_8bit optimizer, with a cosine learning rate scheduler and a learning rate of 5e-6. Additionally, we employ gradient checkpointing, FlashAttention-2 [10], and mixed precision for efficient training. Warm-up steps are set at 10 and weight decay at 0, with checkpoints saved after each epoch. When applying Direct Preference Optimization (DPO) to Janus 7B, the maximum sequence length is adjusted to 2048, micro batch size to 1, epochs to 2, and learning rate to 5e-7, while the rest of the settings remain the same as for instruction tuning. For using ORPO, the alpha is set at 0.1, maximum sequence length at 4096, and epochs at 1, with other settings consistent with those during instruction tuning. For conducting inference on the test set, we use the VLLM library [32]. The maximum number of tokens is set to 4096, with a temperature of 1.0 and a top-p value of 0.9. Additionally, a repetition penalty of 1.03 is specified.
We utilize the OpenRLHF [19] library for implementation. We apply the same training configurations and objective for both reward modeling stages. For each instruction in the training dataset, we concatenate the pair of chosen $y_{c}$ and rejected $y_{r}$ responses prepended with the instruction $x$ as input. The model is trained on each dataset with the reward modeling objective function in Equation 1. The configuration includes maximum sequence length of 2048, micro batch size of 8, batch size of 128. We use the AdamW optimizer with betas of 0.9 and 0.95, learning rate of 9e-6, and weight decay at 0, along with a cosine learning rate scheduler with warmup steps set to 3% of the maximum number of steps. As in E.1, we utilize gradient checkpointing, FlashAttention 2, and mixed precision. DeepSpeed Zero-3 is used for distributed training.
D.2 Computing resources
To train Janus 7B, we utilize four NVIDIA A100 80GB GPUs, and for inference, four NVIDIA RTX A6000 GPUs are employed. Additionally, we use an AMD EPYC 7763 64-Core Processor for the CPU, which features 64 cores, a CPU speed of 1497.674 MHz, and a cache size of 512KB. Instruction tuning, DPO, and ORPO all took approximately  1 day in terms of GPU time, while inference for 1,000 instances took about 10 minutes. The evaluation using the API took approximately  40 minutes.
Appendix E Training details
In this section, we describe the details of training Janus. Section E.1 provides information on Janus 7B, Janus+DPO, and Janus+ORPO. Section E.2 discusses Janus-RM 7B and the methodologies employed, including best-of-n sampling.
E.1 Janus suite training details
E.2 Reward modeling details
We initialize the reward model from Janus 7B trained on the full Multifaceted Collection for 1 epoch, following prior studies [48, 52, 65, 64]. Initializing from the SFT checkpoint ensures that the reward model knows the concept of multifacetedness the pre-trained model learned from a single pass on the entirety of Multifaceted Collection. The classification head for next-token prediction in the pre-trained language model is replaced with a regression head that outputs a scalar reward. We train the reward model in two stages with multifaceted data and helpfulness data, respectively. In the first stage, we use the same pairwise preference data used for training ORPO as the train dataset. In the second stage, we utilize a mix of representative general helpfulness data inspired by [64] and [8], with 72% of HH-RLHF [5], 14% of OASST1 dataset preprocessed for reward modeling [30], and 14% of WebGPT Comparisons [45]. Through extensive experimentations on proximal policy optimization [55] using our trained reward model checkpoints, we empirically observe that continually training on general helpfulness data prevents reward hacking where the model rewards a response directly paraphrasing the given system message.
Appendix F Evaluation details
F.1 Human evaluation
We have employed undergraduate students proficient in English as human evaluators to assess the quality and difficulty of Multifaceted Bench and to compare responses generated by models in a pairwise manner. A total of 15 participants are recruited; six conducted only stage 1, another six participated solely in stage 2, and three took part in both stages. Each task is assigned to nine human evaluators. We utilize Label Studio as the platform for conducting our evaluations.
Stage 1: Multifaceted Bench quality evaluation
We initially divide the 315 unique user instructions found in Multifaceted Bench into nine groups of 35 each. For each user instruction, we match three different system messages with three corresponding reference answers and three evaluation rubrics. Using this setup, we have developed questions to verify quality. The first question asks evaluators to determine if the reference answer adequately responds to both the user instruction and system message, with response options being yes, no, or maybe. The second question requires evaluators to assess whether the provided score rubric adequately addresses the system message, with options of yes or no. (Figure 6) Thus, two questions are assigned for each set of three system message-reference answer-score rubric triples, resulting in six questions per user instruction. Ultimately, each evaluator is assigned 35 user instructions, totaling 210 questions to be assessed. In Figure 10, the human annotation interface for verifying the quality of Multifaceted Bench is depicted.
Figure 6 displays the outcomes of this evaluation. Initially, evaluators found 70% of reference answers to be of high quality, 23.7% to be compliant with user instructions but not system messages, and 6.3% to be of poor quality. Regarding the scoring rubric, 77.2% is considered well-constructed, but 22.8% did not meet the evaluative criteria. This indicates that while Multifaceted Bench exhibits some inconsistencies, it generally maintains good quality, akin to other synthetic datasets. To enhance its utility as a reliable test set, we excluded 24 ‘bad’ samples from a total of 945, leaving 921 samples in the final Multifaceted Bench.
Figure 6: Quality check by human evaluator on Multifaceted Bench. "Yes" indicates good and "No" indicates bad.
Figure 7: Difficulty check of final Multifaceted Bench instances by human evaluator
Stage 2: Model responses evaluation
We conduct human evaluations on Multifaceted Bench by measuring the win-rate directly from human judges for model predictions. A total of three pairs of model comparisons are conducted: Janus vs. Mistral 7B Instruct v0.2, Janus vs. GPT-3.5-Turbo, and Janus vs. GPT-4. Each question consists of two sub-questions. The first sub-question asks judges to assess the overall difficulty of the problem, considering both the instruction and the system message, with options being easy, intermediate, or hard (Figure 7). The second sub-question involves evaluating two model responses based on the score rubric, instruction, system message, and the reference answer, and deciding which response is better, if both are good, or if both are poor.
Initially, we randomly sample one out of three possible system message-reference answer-score rubric sets matched to each unique instruction within Multifaceted Bench, creating 315 sets. These sets are then divided into three groups, with each group receiving 105 sets. We assign the aforementioned pairs of responses to each set within the groups, resulting in a total of 315 sets per group. Consequently, the total number of problems to be evaluated by human judges is 945, and we distribute these among nine evaluators, with each evaluator responding to 105 problems.
Figure 8: Length distribution of responses generated by LLMs and reference answer on Multifaceted Bench
Figure 9: Average / Maximum ROUGE-L scores between pairs of different multifaceted responses per instruction on Multifaceted Bench. The ROUGE-L score is calculated between responses generated by the model for three different system messages given one user instruction.
F.2 LLM-as-a-Judge
In this study, we utilize a proprietary LLM to measure the performance of our model. To evaluate performance within Multifaceted Bench, we provide the evaluator LLM with an instruction, model response, score rubric, and reference answer, and ask it to rate each response on a scale from 1 to 5. We have developed a score rubric for each of the four preference dimensions reflected in the system prompt to assess how well the model adheres to these dimensions. During the evaluation, we input each of the four score rubrics sequentially into the evaluator to perform four assessments for a single model response. Subsequently, the scores generated by the evaluator LLM are averaged to determine the final score for that problem, and the average score across all 921 items is calculated to establish the model’s overall performance score. Furthermore, to ensure the reliability of our LLM-as-a-Judge scoring method, the final scores presented in the tables are averaged from three evaluations per model. The evaluator LLM used in this study is GPT-4-Turbo-0125 with a temperature setting of 1.0. For the evaluation prompts used, please refer to figure J.
F.3 Helpfulness evaluation
To measure helpfulness, we employ three benchmarks: AlpacaEval 2.0, MT-Bench, and Arena Hard Auto v0.1. AlpacaEval 2.0 assesses the Win Rate, specifically how often each model’s generated responses outperform those from the official baseline model, GPT-4-Turbo. A key metric here is the Length Controlled Win Rates (LC Win Rate), which adjusts for length bias. The baseline model in AlpacaEval 2.0 is consistently GPT-4-Turbo. In MT-Bench, evaluator LLMs rate each model’s generated responses on an absolute scale from 0 to 10, using the GPT-4-Turbo model as the evaluator. Finally, Arena Hard Auto v0.1 measures model performance by comparing the win rate against responses generated by GPT-4-Turbo, scoring between 0 and 100 based on this comparison.
F.4 Harmlessness evaluation
To measure harmlessness, we use the RealToxicityPrompts dataset, which evaluates model performance in three aspects: toxicity, fluency, and diversity employing the Perspective API.. For toxicity, we measure two metrics: average maximum toxicity and toxicity probability. To do this, the large language model (LLM) generates 25 responses for each sample from the 10,000-sample test set of RealToxicityPrompts. The highest toxicity among these 25 responses is considered the maximum toxicity for that sample. We then calculate the average of these maximum values to determine the average maximum toxicity. If at least one of the 25 responses is toxic, that sample is counted as toxic, and we calculate the proportion of such cases to determine toxicity probability. Fluency is assessed using the output perplexity of the GPT-2 XL model. Lastly, diversity is measured by the number of unique n-grams, normalized by the text length.
Table 8: System message ablation on Multifaceted Bench. The average score ranges from 1 to 5. If the system message training method is marked as ‘-’, it indicates that the model was not trained using Multifaceted Collection. If the system message given at test time is ‘-’, it means that no system message is provided during the test.
Models System message training method System message given in test time Average
Mistral 7B v0.2 - - 1.012
multifaceted 2.227
Mistral 7B Instruct v0.2 - - 2.778
multifaceted 3.934
GPT-4-Turbo-0125 - - 4.061
multifaceted 4.352
Janus SFT - 3.877
multifaceted 4.171
ORPO - 3.886
multifaceted 4.178
DPO - 3.957
multifaceted 4.242
@@ F.4 Harmlessness evaluation
Table 10: Multifacetedness ablation. ‘-’ indicates that the data was not used for training. MF represents scores ranging from 1 to 5, as measured in the Multifaceted Bench. MT-Bench denotes scores between 1 and 10, as assessed in MT-Bench.
Models System message Response Average
MF MT-Bench
Mistral 7B Instruct v0.2 - - 3.934 7.232
GPT-4-0613 - - 4.271 9.200
Janus - helpful 3.880 7.411
default helpful 3.826 7.532
- multifaceted 4.013 7.612
multifaceted multifaceted 4.171 7.739
Janus* 7B 4.41 4.02 4.36 4.08 4.03 4.14
+ Best-of-64 4.41 4.26 4.42 4.17 4.21 4.28
Janus+DPO 7B 4.45 4.13 4.43 4.21 4.17 4.24
+ Best-of-64 4.49 4.24 4.48 4.20 4.28 4.31
Appendix G Additional analyses
G.1 Qualitative analysis
In our qualitative analysis of responses generated by Janus for Multifaceted Bench, we first compare the length distribution of responses from Janus, other baseline models, and the reference answers. As illustrated in Figure 8, responses from Janus are, on average, longer than those from Mistral 7B Instruct v0.2, GPT-3.5 Turbo-0125, and GPT-4-Turbo-0125. Notably, the length distribution of responses from Janus 7B closely matches that of the reference answers. Furthermore, Janus 7B not only aligns well with the length distribution but also generates qualitatively superior responses compared to other models. For example, as shown in Table LABEL:tab:example_alternate_expertise, when given system message that require a straightforward, logic-based approach, Janus 7B accurately follows this directive and constructs logical responses. In contrast, while the responses from GPT-4 are not entirely incorrect, they still fail to adhere closely to the user’s preferences. These results demonstrate that Janus not only performs well on benchmarks but also effectively reflects individual human preferences in the quality of its generated responses. For additional examples, please refer to Section K.
Appendix H Limitations
It is essential to recognize that the data may still face limitations tied to the quality of synthetic training data. Direct human evaluation is unfeasible, and the data may include samples of lower quality. Nevertheless, the performance of Janus indicates that the majority of the data is of good quality. Furthermore, Janus has been demonstrated to be relatively harmless compared to other LLMs. However, potential ethical concerns related to privacy, consent, or misuse could still emerge. Additionally, during the creation of Multifaceted Collection, preferences were collected with consideration of the harmlessness dimension. Consequently, Janus, trained with this data, exhibits lower toxicity levels compared to baseline toxicity measurement benchmarks, thereby reducing the likelihood of generating harmful responses. Despite these precautions, there remains a risk of generating harmful responses in the face of unforeseen malicious attacks, such as jailbreaking.
Appendix I Broader impact
Enhanced user satisfaction and engagement Janus responses according to individual preferences could significantly enhance user satisfaction by making interactions more relevant and engaging. This could lead to wider adoption of LLMs in consumer applications, such as personalized learning, and customer service.
Better inclusion and representation By addressing diverse and nuanced human preferences, our approach could lead to more inclusive technology that better represents and serves users from different backgrounds and with different values. This could help mitigate biases commonly seen in technology that is trained on more homogeneous or mainstream data.
Reduction in retraining costs and resource use Our approach can reduce the need for frequent retraining of models to accommodate new user preferences. This could lower computational costs and energy consumption associated with training large models, contributing to more sustainable AI practices.
Improved data privacy Personalizing without the need to continually collect and store detailed real human user data could have implications for privacy. By reducing the amount of data needed and processing it in a way that doesn’t require detailed personal histories, there’s potential for enhancing user trust and complying with strict data privacy regulations.

### v2

@@ Appendix D Details of Multifaceted Collection construction
Nectar anthropic-hh 6,037
lmsys-chat-1m 3,119
sharegpt 1,994
ultrachat 783
evol_instruct 775
flan_v2_niv2 583
flan_v2_cot 253
false_qa 176
flan_v2_p3 171
truthful_qa 54
flan_v2_flan2021 33
@@
Seed value creation
We hand-crafted 4 dimensions, 18 sub-dimensions, and 107 value descriptions. The full list keywords of our seed data are in Table 13.
Table 13: Keyword of 107 seed values
Dimension Subdimension Value keywords
@@
Appendix E Analysis of Multifaceted Collection
Diversity of preferences
To measure the diversity of preferences in Multifaceted Collection, we computed ROUGE-L999pypi.org/project/rouge-score/ (F1 score) similarities between every possible pair of preference descriptions assigned to each instruction. Given that there are three preference sets per instruction and four dimensions defined in each preference set, we computed the metric along the four dimensions, i.e., $\binom{3}{2}\times 4=12$ scores are obtained per instruction. As depicted in Figure 9, the average ROUGE-L score across all dimensions is approximately 0.21, with the maximum score reaching 0.25. Compared to results from previous studies creating synthetic datasets [20, 81], there is significant diversity among the preferences. It is observed they do not overlap substantially and are well-distributed within Multifaceted Collection. Example subdimensions and keywords of preference descriptions are displayed in Table 14.
Quality of system messages
We develop two criteria that a proper system message should adhere to: 1) relevance and specificity and 2) coherence and naturalness. We create a score rubric indicating a scale of 1 to 5 for each criterion. Inspired by recent works that use LLM-as-a-Judge to assess the process of training data synthesis [89, 85], we use LLaMA 3.1 8B Instruct [45] to score a random 20k subset of system message-user instruction pairs. Results show an average of 3.74 on relevance and specificity and 4.01 on coherence and naturalness, with 68.8% and 85.6% instances at or above score 4, respectively. This demonstrates the quality of verbalized preferences, potentially revealing why our model is effectively steerable for pluralistic alignment.
Safety
To check the presence of unsafe content in our synthetic dataset, we evaluate the dataset’s system messages, user prompts, and gold responses using a content safety classifier, Llama Guard 3 8B [45]. 99.2% of the 196,998 instances are classified as safe.
@@ Appendix E Analysis of Multifaceted Collection
Appendix F Ablation data: default system messages and generally helpful responses
@@
We refer to system messages set as default in popular LLMs as default system messages. Table 15 shows a list of default system messages that contain no model-specific instructions and are sourced from FastChat101010github.com/lm-sys/FastChat/blob/main/fastchat/conversation.py.
To generate generally helpful responses used for training in Table 7, we use same instructions in Multifaceted Collection but with every three multifaceted system messages per instruction replaced with three default system messages randomly sampled from Table 7. Then, GPT-4-Turbo-0125 generate gold responses for the pair of default system message and instruction similar to the Multifaceted Collection construction process.
Appendix G Details of Multifaceted Bench construction
Instruction sampling
We sample instructions from five benchmarks and exclude instructions that are either used in Multifaceted Collection or overlapped across the benchmarks, and then sample from the remaining messages to collect 315 unique instructions.
Preference set, system message, gold response generation
We use GPT-4 Turbo-0125 to generate synthetic multifaceted system messages and their aligned reference answers, similar to how we curate the Multifaceted Collection. For each instruction, we generate three system messages and one reference answer, resulting in a total of 945 sets. In this process, we generate system messages with unseen preferences that are not included in the Multifaceted Collection, thereby reducing the possibility of data contamination.
Score rubric generation
In line with previous works [92, 30, 36, 31] indicating that using a score rubric helps in conducting consistent evaluations, we also develop a score rubric to assess the reference answers. The score rubric evaluates how well the model’s response adheres to each of the four preference dimensions that comprise the system message. We create one rubric per dimension, resulting in four rubrics for each response. Each rubric includes an explanation of the dimension and criteria for scoring from 1 to 5 based on how well the model follows it. The generation prompt for score rubrics is in Appendix M.
Instance quality annotation and filtering
We employ nine human annotators to assess the quality of the test dataset. See Appendix K for annotator details. We initially divide the 315 unique user instructions into 9 groups of 35 each. For each user instruction, we match three different system messages with three corresponding reference answers and three evaluation rubrics. Annotators are tasked to respond to two questions. The first question asks to determine if the reference answer adequately responds to both the user instruction and system message, with response options being “yes”, “no”, or “maybe”. The second question requires to assess whether the provided score rubric adequately addresses the system message, with options of “yes” or “no”. Thus, two questions are assigned for each set of three (system message, reference answer, score rubric) triples, resulting in six questions per user instruction. Ultimately, each evaluator is assigned 35 user instructions, totaling 210 questions to be assessed. The annotation interface is displayed in Figure 13.
Figure 10 shows annotation results. Initially, annotators found 70% of reference answers to be of high quality, 23.7% to be compliant with user instructions but not system messages, and 6.3% to be of poor quality. Regarding the scoring rubric, 77.2% is considered well-constructed, but 22.8% did not meet the evaluative criteria. This indicates that while Multifaceted Bench exhibits some inconsistencies, it generally maintains good quality, akin to other synthetic datasets. To enhance its utility as a reliable test set, we excluded 24 “bad” samples from a total of 945, leaving 921 samples in the final Multifaceted Bench.
Figure 10: Quality check by human evaluator on Multifaceted Bench. “Yes” indicates good and “No” indicates bad.
Appendix H Analysis of Multifaceted Bench
Figure 12: Difficulty distribution of final Multifaceted Bench instances by human evaluator
Train-test similarity
To verify whether Multifaceted Bench indeed contains unseen preferences not present in Multifaceted Collection, we measure the overlap between the system prompts of the two datasets by calculating their Rouge-L scores. As shown in Figure 12, the average ROUGE-L score is average 0.17 and maximum 0.28 among three system messages per instruction, indicating a sufficiently low overlap. Additionally, this score is lower than the similarities among training data as depicted in figure 9. This low level of train-test similarity, compared to previous studies [81, 30], confirms that Multifaceted Bench successfully represents unseen tasks.
Difficulty of instances
Figure 12 shows the difficulty of Multifaceted Bench instances in 3 levels measured by human annotators. See Appendix K and Appendix L.1 for the detailed procedure.
Appendix I Computing resources
To train Janus 7B, we utilize four NVIDIA A100 80GB GPUs, and for inference, four NVIDIA RTX A6000 GPUs are employed. Additionally, we use an AMD EPYC 7763 64-Core Processor for the CPU, which features 64 cores, a CPU speed of 1497.674 MHz, and a cache size of 512KB. Instruction tuning, DPO, and ORPO all took approximately  1 day in terms of GPU time, while inference for 1,000 instances took about 10 minutes. Evaluation using the API took approximately 4̃0 minutes.
Appendix J Training and inference details
J.1 Instruction tuning and preference tuning
Data composition
Hyperparameters and implementation
To train Janus 7B, we utilize the axolotl library111111github.com/OpenAccess-AI-Collective/axolotl. For instruction tuning, the configuration includes a maximum sequence length of 8192, gradient accumulation steps of 4, a micro batch size of 2, and four epochs. We use the adamw_bnb_8bit optimizer, with a cosine learning rate scheduler and a learning rate of 5e-6. Additionally, we employ gradient checkpointing, FlashAttention-2 [10], and mixed precision for efficient training. Warm-up steps are set at 10 and weight decay at 0, with checkpoints saved after each epoch. When applying Direct Preference Optimization (DPO) to Janus 7B, the maximum sequence length is adjusted to 2048, micro batch size to 1, epochs to 2, and learning rate to 5e-7, while the rest of the settings remain the same as for instruction tuning. For using ORPO, the alpha is set at 0.1, maximum sequence length at 4096, and epochs at 1, with other settings consistent with those during instruction tuning.
J.2 Reward modeling
Model selection
We initialize the reward model from Janus 7B trained on the full Multifaceted Collection for 1 epoch, following prior studies [56, 61, 78, 77]. Initializing from the SFT checkpoint ensures that the reward model knows the concept of multifacetedness the pre-trained model learned from a single pass on the entirety of Multifaceted Collection. The classification head for next-token prediction in the pre-trained language model is replaced with a regression head that outputs a scalar reward.
Data composition
We train the reward model in two stages with multifaceted data and helpfulness data, respectively. In the first stage, we use the same pairwise preference data used for training ORPO as the train dataset. In the second stage, we utilize a mix of representative general helpfulness data inspired by [77] and [8], with 72% of HH-RLHF [5], 14% of OASST1 dataset preprocessed for reward modeling [33], and 14% of WebGPT Comparisons [52]. Through extensive experimentations on proximal policy optimization [66] using our trained reward model checkpoints, we empirically observe that continually training on general helpfulness data prevents reward hacking where the model rewards a response directly paraphrasing the given system message.
Hyperparameters and implementation
We use the OpenRLHF [22] library for training. The following training configurations and objective are applied to both reward modeling stages. For each instruction in the training dataset, we concatenate the pair of chosen $y_{c}$ and rejected $y_{r}$ responses prepended with the instruction $x$ as input. The model is trained on each dataset with the reward modeling objective function in Equation 1. The configuration includes maximum sequence length of 2048, micro batch size of 8, batch size of 128. We use the AdamW optimizer with betas of 0.9 and 0.95, learning rate of 9e-6, and weight decay at 0, along with a cosine learning rate scheduler with warmup steps set to 3% of the maximum number of steps. As in Appendix J.1, we utilize gradient checkpointing, FlashAttention 2, and mixed precision. DeepSpeed Zero-3 is used for distributed training.
J.3 Inference
Hyperparameters and implementation
For inference on test datasets, we use the VLLM library [35]. The maximum number of tokens is set to 4096, with a temperature of 1.0 and a top-p value of 0.9. Additionally, a repetition penalty of 1.03 is specified.
Appendix K Human annotators information
Human annotators are employed in order to assess the quality and difficulty of Multifaceted Bench (Stage 1) and to compare responses generated by models in a pairwise manner (Stage 2).
We employ 14 English-proficient Korean undergraduate students aged 20-24 from various majors, consisting of 6 males and 8 females. 6 participated only in Stage 1, another 5 participated solely in Stage 2, and 3 took part in both stages. Each task is assigned to nine human workers.
Each stage lasted two hours. We pay ₩40,000 per person for completing a single annotation stage.
The annotation interface and instructions for Stage 1 is in Figure 13 and the interface for Stage 2 is in Figure 14.
Appendix L Evaluation details
L.1 Multifacetedness evaluation
Human evaluation
We measure win rate of model responses to Multifaceted Bench by employing eight human evaluators. See Appendix K for worker details. A total of three pairs of model comparisons are conducted: Janus vs. Mistral 7B Instruct v0.2, Janus vs. GPT-3.5-Turbo, and Janus vs. GPT-4. Each question consists of two sub-questions. The first sub-question asks judges to assess the overall difficulty of the problem, considering both the instruction and the system message, with options being easy, intermediate, or hard. The second sub-question involves evaluating two model responses based on the score rubric, instruction, system message, and the reference answer, and deciding which response is better, if both are good, or if both are poor. The annotation interface is in Figure 14. Initially, we randomly sample one out of three possible system message-reference answer-score rubric sets matched to each unique instruction within Multifaceted Bench, creating 315 sets. These sets are then divided into 3 groups, with each group receiving 105 sets. We assign the aforementioned pairs of responses to each set within the groups, resulting in a total of 315 sets per group. Consequently, the total number of problems to be evaluated by human judges is 945, and we distribute these among nine evaluators, with each evaluator responding to 105 problems.
LLM-as-a-Judge
We additionally utilize a proprietary LLM to measure the performance of our model. To evaluate performance within Multifaceted Bench, we provide the evaluator LLM with an instruction, model response, score rubric, and reference answer, and ask it to rate each response on a scale from 1 to 5. We have developed a score rubric for each of the four preference dimensions reflected in the system prompt to assess how well the model adheres to these dimensions. During the evaluation, we input each of the four score rubrics sequentially into the evaluator to perform four assessments for a single model response. Subsequently, the scores generated by the evaluator LLM are averaged to determine the final score for that problem, and the average score across all 921 items is calculated to establish the model’s overall performance score. Furthermore, to ensure the reliability of our LLM-as-a-Judge scoring method, the final scores presented in the tables are averaged from three evaluations per model. The evaluator LLM used in this study is GPT-4-Turbo-0125 with a temperature setting of 1.0. For the evaluation prompts used, please refer to Appendix M.
L.2 Helpfulness evaluation
AlpacaEval 2.0
The Win Rate is measured, specifically how often each model’s generated responses outperform those from the official baseline model, GPT-4-Turbo. A key metric here is the Length Controlled Win Rates (LC Win Rate), which adjusts for length bias. The baseline model in AlpacaEval 2.0 is consistently GPT-4-Turbo.
MT-Bench
Evaluator LLMs rate each model’s generated responses on an absolute scale from 0 to 10, using the GPT-4-Turbo model as the evaluator.
Arena Hard Auto v0.1
Model performance is measured by comparing the win rate against responses generated by GPT-4-Turbo, scoring between 0 and 100 based on this comparison.
L.3 Harmlessness evaluation
RealToxicityPrompts
The model performance is evaluated in three aspects: toxicity, fluency, and diversity employing the Perspective API. For toxicity, we measure two metrics: average maximum toxicity and toxicity probability. To do this, the large language model (LLM) generates 25 responses for each sample from the 10,000-sample test set of RealToxicityPrompts. The highest toxicity among these 25 responses is considered the maximum toxicity for that sample. We then calculate the average of these maximum values to determine the average maximum toxicity. If at least one of the 25 responses is toxic, that sample is counted as toxic, and we calculate the proportion of such cases to determine toxicity probability. Fluency is assessed using the output perplexity of the GPT-2 XL model. Lastly, diversity is measured by the number of unique n-grams, normalized by the text length.
For the three social bias benchmarks below, we run all of them in zero-shot using a specific branch of the Language Model Evaluation Harness [15]121212github.com/EleutherAI/lm-evaluation-harness/tree/winogender. We include Gemma 2 9B IT [76] as a baseline as it is a state-of-the-art similar-sized model reported to have been extensively tested on bias benchmarks.
Winogender
We calculate the accuracy of pronoun prediction for a particular gender, namely male, female, and neutral. We also examine accuracy in ‘gotcha’ scenarios in the dataset. These refer to a subset of the dataset where the gender of the pronoun referring to an occupation does not match U.S. statistics on that occupation’s majority gender, i.e., they challenge the model’s reliance on stereotypes more.
CrowS-Pairs
The metrics are: (i) the average absolute difference in log-likelihoods between stereotypical and non-stereotypical sentence pairs and (ii) the rate of how often the model considers stereotypical sentences more likely than non-stereotypical ones.
BBQ
Model bias is assessed in two scenarios: (i) whether a model falls back on stereotypes when given limited context (“Ambiguous”) and (ii) whether the model’s biases override a correct choice even when provided with clear context (“Disambiguated”). The bias score quantifies the degree to which a model’s responses align with societal stereotypes or biases, ranging from -100% (completely against the bias) to 100% (fully aligned with the bias), with 0% indicating no measured bias.